# Smooth Sailing: Lipschitz-Driven Uncertainty Quantification for Spatial Association

### v2

@@ Appendix C Simulations Additional Details
C.1 Reported Uncertainty in Simulation Experiments
In Fig. 1, we provide error bars for the empirical coverage as well as a point estimate. The upper side of the confidence interval indicates the largest value of the (true) coverage such that with probability $97.5\%$, the empirical coverage would be less than or equal to the observed value. Conversely, the lower edge of the confidence interval indicates the smallest value of the (true) coverage such that with probability $97.5\%$, the empirical coverage would be greater than or equal to the observed value. This is therefore a (conservative) $95\%$ confidence interval for the true coverage.
@@ C.1 Reported Uncertainty in Simulation Experiments
We also provide ±$\pm 2$ standard deviation error bars on the confidence interval width. As the confidence interval width is not necessarily normal this may not be a $95\%$ confidence interval for the confidence interval width. But is meant as an indicator of spread of confidence interval widths each method obtains.
C.2 Data Generation for the Single Covariate Experiment
@@
We show example datasets used in the simulation experiment with a single covariate and the simulation experiment in which we investigated the impact of Lipschitz constant on confidence interval width and coverage in Fig. 4. The left most two panels show the distribution of source (blue) and target (orange) locations for two values of the shift parameter. Large positive values of the shift parameter lead to target distributions clustered to the top right, values close to zero lead to the target locations being approximately uniformly distributed and large negative values lead to target locations clustered to the bottom left. The third panel from the left in Fig. 4 shows the covariate plotted as a function of spatial location,
@@
This obtains a maximum of $2\sqrt{2}$ on $[-1,1]^{2}$, and so this is the Lipschitz constant of $f(S)=\mathbb{E}[Y|S]$.
C.3 Data Generation for the Three Covariate Experiment
In Fig. 5 we show data generated for the three covariate shift experiment. Source and target locations are generated as in the one covariate simulation, but with $N=10{,}000$ ($M=100$ is still used). These are not shown in Fig. 5.
The covariates are, from left to right in Fig. 5
@@ C.3 Data Generation for the Three Covariate Experiment
The conditional expectation of the response (right most panel in Fig. 5) is
@@
Figure 5: The first 3 plots from the left show the covariate surfaces, while the fourth shows the expected response at each spatial location for the second simulated experiment. The source and target locations (not shown) are the same as in Fig. 4, though with $N=10{,}000$.
Appendix D Tree Cover Experiment Additional Details
@@
D.1 Tree Cover Data
@@
Figure 6 provides a visual overview of both the tree cover and covariates. As a preprocessing step, we convert the (latitude, longitude) coordinates of each data point into radians. This allows us to use the Haversine formula to compute distances in kilometers for the Wasserstein-1 cost and the nearest neighbor weighting procedure.
@@
Figure 6 provides a visual overview of the distribution for both the tree cover and covariates.
@@
D.2 Source and Target Data Split and Spatial Preprocessing
We define our target region to be the Western portion of the Continental United States. In particular, we consider locations that have latitudes between $25^{\circ}$ and $50^{\circ}$ and longitudes between $-125^{\circ}$ and $-110^{\circ}$. Within spatial locations in this defined region, we pick $50\%$ of all spatial locations — totaling 133 sites — as target data.
@@ D.2 Source and Target Data Split and Spatial Preprocessing
Fig. 7 visually represents the spatial distribution of the source and target locations for one representative random seed.
@@
D.3 Estimating the Ground Truth Parameters to Evaluate Coverage
@@
where $\hat{\theta}_{p}^{\star}=e_{p}^{\mathrm{T}}(X^{\star\mathrm{T}}X^{\star})^{-1}X^{\star\mathrm{T}}Y^{\star}$ is the estimated parameter using the target data (which our method and baselines don’t have access to) and $\hat{\theta}_{p}$ is the estimated parameter we compute with each method using the source responses. If this confidence interval contains 00, we count the method as having covered the true parameter, while if it doesn’t we count the method as not having covered the true parameter. We estimate the variance of $\hat{\theta}_{p}^{\star}$ using the model-trusting standard errors $\hat{\sigma}^{2}=\frac{1}{N-P}\sum_{n=1}^{N}r_{n}^{2}$ where $r_{n}$ are the residuals of the model fit on the training data. We expect this to inflate our estimate of the standard variance of the target OLS estimate if the response surface is nonlinear, as the residuals will be larger due to bias. By possibly overestimating and incorporating this sampling variability into confidence intervals, we expect the calculated coverages to overestimate the true coverages. The resulting coverages are shown in the top row of Fig. 3.
In Fig. 8 we also report confidence intervals by calculating the proportion of times that $\hat{\theta}_{p}^{\star}$ is contained in each confidence interval. $\hat{\theta}_{p}$ is an unbiased estimate for $\theta^{\star}_{\text{OLS},p}$ whether or not the model is well-specified. However, we might expect that the coverages reported with this approach underestimate the actual coverage of each method’s confidence intervals due to not accounting for sampling variability in $\hat{\theta}_{p}^{\star}$.
@@
D.4 Experiment with Varying Lipschitz Constant
In this section, we provide an additional experiment with these real data where the focus is to assess how varying the underlying Lipschitz constant in Assumption 4 changes coverage and interval width. For this experiment we consider 9 different values for the Lipschitz constant, $L\in\{0.001,0.005,0.01,0.02,0.05,0.1,0.2,0.5,1\}$. This range of values is a reasonable range of values that a practitioner with domain knowledge in the tree cover field might be willing to specify. Indeed, the smallest Lipschitz constant considered, $0.001$, corresponds to assuming that to have a $1\%$ increase in tree cover we need to move $1/0.001=1000$ kilometers. This is quite an extreme value, but in some parts of arid regions such as New Mexico and Arizona it can be true. On the other hand, the largest Lipschitz constant considered, $1$, corresponds to assuming that to have $1\%$ increase in tree cover we need to move $1$ kilometer. This is also extreme, but in regions in the US where elevation changes are very pronounced (e.g. in the Colorado Rockies), tree cover can vary sharply over very short distances.
In Fig. 9, we present the results of this experiment. We find that varying the Lipschitz constant within this range does not significantly impact coverage. Specifically, for slope and elevation, coverage remains consistent across all constants except $L=0.001$. For aridity index, coverage is low for $L\leq 0.1$ but exceeds $95\%$ for $L\geq 0.2$. Meanwhile, confidence intervals become noticeably wider for $L>0.5$ while remaining relatively stable for smaller values.
@@ D.4 Experiment with Varying Lipschitz Constant
D.5 Additional Experiment: Target South-East US
In this experiment, we define our target region in the Southeastern portion of CONUS at locations with latitude in the range (25, 38) and longitude in the range (-100, -75). Out of all spatial points in this region, $50\%$ — totaling 118 sites — are designated as target data. Next, we select the source data by taking a uniform random sample of $20\%$ of the remaining spatial locations, repeated over 250 random seeds to assess coverage performance. Each seed yields 173 source locations. Fig. 11 illustrates the spatial split between source and target data for a representative seed. As before, as a preprocessing step we convert the (latitude, longitude) coordinates of each data point into radians.
@@ Results.
We report the results for the confidence interval coverage and width in Fig. 12. As before, our method consistently achieves or exceeds 95% nominal coverage for all parameters. All competing methods but KDEIW here achieves $95\%$ nominal coverage (or close to $95\%$ nominal coverage) for the aridity index and slope parameter when considering the coverage for the $\hat{\theta_{p}^{\star}}-\hat{\theta}_{p}$ (top row). All competing methods fall short of the nominal threshold for the elevation parameter. The wider intervals produced by our method (last row) reflect the trade-off between achieving reliable coverage and maintaining narrower intervals.
@@
Finally, we report also for this experiment results when varying the assumed Lipschitz constant. As explained in Section D.4, we consider 9 different values for the Lipschitz constant, $L\in\{0.001,0.005,0.01,0.02,0.05,0.1,0.2,0.5,1\}$. We report the results in Fig. 14. As before, we find that varying the Lipschitz constant within this range does not significantly impact coverage. Specifically, here we see that for aridity index and elevation, coverage remains consistent across all constants above $L=0.01$. For slope, $95\%$ nominal coverage is achieved for $L\geq 0.05$. And — as before — confidence intervals become noticeably wider for $L>0.5$ while remaining relatively stable for smaller values.
@@
Appendix E Proofs
E.1 Derivation of Target-Conditional Ordinary Least Squares
@@
  • The second term, $-2\mathbb{E}\left[\sum_{m=1}^{M}Y^{\star}_{m}\theta^{\top}X^{\star}_{m}\Big{|}S^{\star}_{m}\right]$, can be rewritten using the linearity of expectation, the fact that $\theta$ is not random, and Assumption 1:
@@
  • The third term, $\mathbb{E}\left[\sum_{m=1}^{M}(\theta^{\top}X^{\star}_{m})^{2}\Big{|}S^{\star}_{m}\right]$, is non-random by Assumption 1 and can be rewritten as:
@@
as in Eq. 4.
E.2 Proof of Theorem 7
In the main text, we stated our confidence interval for the $1$NN choice of $\Psi$. We now state version for general non-negative matrices with columns summing to $1$.
@@ E.2 Proof of Theorem 7
Suppose $(S^{\star}_{m},X^{\star}_{m},Y^{\star}_{m})_{m=1}^{M}$ and $(S_{n},X_{n},Y_{n})_{n=1}^{N}$ satisfy Assumptions 1, 4, 3 and 2. Define $w$ as in Algorithm 1. For any $\Psi\in\mathbb{R}^{M\times N}$, a matrix with non-negative entries with columns summing to $1$, define $v^{\Psi}=w\Psi$. As in the main text, take $\hat{\theta}_{p}^{\Psi}=e_{p}^{\mathrm{T}}\left(X^{\star\mathrm{T}}X^{\star}\right)^{-1}X^{\star\mathrm{T}}\Psi Y$. Define $c=\sigma\|v^{\Psi}\|_{2}$, with $\sigma^{2}$ the variance of the additive noise from Assumption 2. Define the (random) interval $I^{\Psi}$ as in Algorithm 1 with known $\sigma^{2}$. Then with probability at least $1-\alpha$, $\theta^{\star}_{\text{OLS},p}\in I^{\Psi}$. That is, $I^{\Psi}$ has coverage (conditional on the test locations) at least $1-\alpha$.
@@
for $w:=e_{p}^{\mathrm{T}}\left(X^{\star\mathrm{T}}X^{\star}\right)^{-1}X^{\star\mathrm{T}}\in\mathbb{R}^{M}$ and $v^{\Psi}:=w\Psi\in\mathbb{R}^{N}$. By Assumption 2, the righthand side of Eq. 47 can be written
@@
We used the $\epsilon_{n}$ are independent and identically distributed with variance $\sigma^{2}$ (Assumption 2).
@@
To bound the bias $b$, we use Assumption 4 to write
@@
We can therefore apply Lemma 6 to complete the proof. ∎
E.3 Computing an Upper Bound on the Bias of Our Estimand with Wasserstein-$1$ Distance
First of all observe that if $\sum_{m=1}^{M}w_{m}=1,\sum_{n=1}^{N}v_{n}^{\Psi}=1$, $v_{n}^{\Psi}\geq 0$ for $1\leq n\leq N$, and $w_{m}\geq 0$ for $1\leq m\leq M$ then the supremum in Eq. 7 would be equal to a Wasserstein-$1$ distance by Kantorovich-Rubinstein duality [65, Theorem 5.10, Case 5.16].
Next, consider what happens if $\sum_{m=1}^{M}w_{m}-\sum_{n=1}^{N}v_{n}^{\Psi}\neq 0$. We show that the right-hand side of Eq. 7 is infinite. Assume by contradiction that there exists a $C>0$ that upper bounds this supremum. Because $\sum_{m=1}^{M}w_{m}-\sum_{n=1}^{N}v_{n}^{\Psi}\neq 0$ and all constant functions are $L$-Lipschitz, for any $\gamma>0$, taking, for all $S$, $g(S)=G=\frac{C+\gamma}{\sum_{m=1}^{M}w_{m}-\sum_{n=1}^{N}v_{n}^{\Psi}}$

### v3

@@ Appendix D Simulations Additional Details
D.1 Reported Uncertainty in Simulation Experiments
In Fig.˜1, we provide error bars for the empirical coverage as well as a point estimate. The upper side of the confidence interval indicates the largest value of the (true) coverage such that with probability $97.5\%$, the empirical coverage would be less than or equal to the observed value. Conversely, the lower edge of the confidence interval indicates the smallest value of the (true) coverage such that with probability $97.5\%$, the empirical coverage would be greater than or equal to the observed value. This is therefore a (conservative) $95\%$ confidence interval for the true coverage.
@@ D.1 Reported Uncertainty in Simulation Experiments
We also provide $\pm 2$ standard deviation error bars on the confidence interval width. As the confidence interval width is not necessarily normal this may not be a $95\%$ confidence interval for the confidence interval width. But is meant as an indicator of spread of confidence interval widths each method obtains.
D.2 Data Generation for the Single Covariate Experiment
@@
We show example datasets used in the simulation experiment with a single covariate and the simulation experiment in which we investigated the impact of Lipschitz constant on confidence interval width and coverage in Fig.˜4. The left most two panels show the distribution of source (blue) and target (orange) locations for two values of the shift parameter. Large positive values of the shift parameter lead to target distributions clustered to the top right, values close to zero lead to the target locations being approximately uniformly distributed and large negative values lead to target locations clustered to the bottom left. The third panel from the left in Fig.˜4 shows the covariate plotted as a function of spatial location,
@@
This obtains a maximum of $2\sqrt{2}$ on $1,1]^{2}$, and so this is the Lipschitz constant of $f(S)=\mathbb{E}[Y|S]$.
D.3 Data Generation for the Three Covariate Experiment
In Fig.˜5 we show data generated for the three covariate shift experiment. Source and target locations are generated as in the one covariate simulation, but with N=10, [^0] N=10{,}000 ($M=100$ is still used). These are not shown in Fig.˜5.
The covariates are, from left to right in Fig.˜5
@@ D.3 Data Generation for the Three Covariate Experiment
The conditional expectation of the response (right most panel in Fig.˜5) is
@@
Figure 5: The first 3 plots from the left show the covariate surfaces, while the fourth shows the expected response at each spatial location for the second simulated experiment. The source and target locations (not shown) are the same as in Fig.˜4, though with N=10, [^0] N=10{,}000.
Appendix E Tree Cover Experiment Additional Details
@@
E.1 Tree Cover Data
@@
Figure˜6 provides a visual overview of both the tree cover and covariates. As a preprocessing step, we convert the (latitude, longitude) coordinates of each data point into radians. This allows us to use the Haversine formula to compute distances in kilometers for the Wasserstein-1 cost and the nearest neighbor weighting procedure.
@@
Figure˜6 provides a visual overview of the distribution for both the tree cover and covariates.
@@
E.2 Source and Target Data Split and Spatial Preprocessing
We define our target region to be the Western portion of the Continental United States. In particular, we consider locations that have latitudes between $25^{\circ}$ and $50^{\circ}$ and longitudes between $125^{\circ}$ and $110^{\circ}$. Within spatial locations in this defined region, we pick $50\%$ of all spatial locations — totaling 133 sites — as target data.
@@ E.2 Source and Target Data Split and Spatial Preprocessing
Fig.˜7 visually represents the spatial distribution of the source and target locations for one representative random seed.
@@
E.3 Estimating the Ground Truth Parameters to Evaluate Coverage
@@
where $\hat{\theta}_{p}^{\star}=e_{p}^{\mathrm{T}}(X^{\star\mathrm{T}}X^{\star})^{-1}X^{\star\mathrm{T}}Y^{\star}$ is the estimated parameter using the target data (which our method and baselines don’t have access to) and $\hat{\theta}_{p}$ is the estimated parameter we compute with each method using the source responses. If this confidence interval contains 0, we count the method as having covered the true parameter, while if it doesn’t we count the method as not having covered the true parameter. We estimate the variance of $\hat{\theta}_{p}^{\star}$ using the model-trusting standard errors $\hat{\sigma}^{2}=\frac{1}{N-P}\sum_{n=1}^{N}r_{n}^{2}$ where $r_{n}$ are the residuals of the model fit on the training data. We expect this to inflate our estimate of the standard variance of the target OLS estimate if the response surface is nonlinear, as the residuals will be larger due to bias. By possibly overestimating and incorporating this sampling variability into confidence intervals, we expect the calculated coverages to overestimate the true coverages. The resulting coverages are shown in the top row of Fig.˜3.
In Fig.˜8 we also report confidence intervals by calculating the proportion of times that $\hat{\theta}_{p}^{\star}$ is contained in each confidence interval. $\hat{\theta}_{p}$ is an unbiased estimate for $\theta^{\star}_{\text{OLS},p}$ whether or not the model is well-specified. However, we might expect that the coverages reported with this approach underestimate the actual coverage of each method’s confidence intervals due to not accounting for sampling variability in $\hat{\theta}_{p}^{\star}$.
@@
E.4 Experiment with Varying Lipschitz Constant
In this section, we provide an additional experiment with these real data where the focus is to assess how varying the underlying Lipschitz constant in Assumption˜4 changes coverage and interval width. For this experiment we consider 9 different values for the Lipschitz constant, $L\in\{0.001,0.005,0.01,0.02,0.05,0.1,0.2,0.5,1\}$. This range of values is a reasonable range of values that a practitioner with domain knowledge in the tree cover field might be willing to specify. Indeed, the smallest Lipschitz constant considered, $0.001$, corresponds to assuming that to have a $1\%$ increase in tree cover we need to move $1/0.001=1000$ kilometers. This is quite an extreme value, but in some parts of arid regions such as New Mexico and Arizona it can be true. On the other hand, the largest Lipschitz constant considered, $1$, corresponds to assuming that to have $1\%$ increase in tree cover we need to move $1$ kilometer. This is also extreme, but in regions in the US where elevation changes are very pronounced (e.g. in the Colorado Rockies), tree cover can vary sharply over very short distances.
In Fig.˜9, we present the results of this experiment. We find that varying the Lipschitz constant within this range does not significantly impact coverage. Specifically, for slope and elevation, coverage remains consistent across all constants except $L=0.001$. For aridity index, coverage is low for $L\leq 0.1$ but exceeds $95\%$ for $L\geq 0.2$. Meanwhile, confidence intervals become noticeably wider for $L>0.5$ while remaining relatively stable for smaller values.
@@ E.4 Experiment with Varying Lipschitz Constant
E.5 Additional Experiment: Target South-East US
In this experiment, we define our target region in the Southeastern portion of CONUS at locations with latitude in the range (25, 38) and longitude in the range (-100, -75). Out of all spatial points in this region, $50\%$ — totaling 118 sites — are designated as target data. Next, we select the source data by taking a uniform random sample of $20\%$ of the remaining spatial locations, repeated over 250 random seeds to assess coverage performance. Each seed yields 173 source locations. Fig.˜11 illustrates the spatial split between source and target data for a representative seed. As before, as a preprocessing step we convert the (latitude, longitude) coordinates of each data point into radians.
@@ Results.
We report the results for the confidence interval coverage and width in Fig.˜12. As before, our method consistently achieves or exceeds 95% nominal coverage for all parameters. All competing methods but KDEIW here achieves $95\%$ nominal coverage (or close to $95\%$ nominal coverage) for the aridity index and slope parameter when considering the coverage for the $\hat{\theta_{p}^{\star}}-\hat{\theta}_{p}$ (top row). All competing methods fall short of the nominal threshold for the elevation parameter. The wider intervals produced by our method (last row) reflect the trade-off between achieving reliable coverage and maintaining narrower intervals.
@@
Finally, we report also for this experiment results when varying the assumed Lipschitz constant. As explained in Section˜E.4, we consider 9 different values for the Lipschitz constant, $L\in\{0.001,0.005,0.01,0.02,0.05,0.1,0.2,0.5,1\}$. We report the results in Fig.˜14. As before, we find that varying the Lipschitz constant within this range does not significantly impact coverage. Specifically, here we see that for aridity index and elevation, coverage remains consistent across all constants above $L=0.01$. For slope, $95\%$ nominal coverage is achieved for $L\geq 0.05$. And — as before — confidence intervals become noticeably wider for $L>0.5$ while remaining relatively stable for smaller values.
@@
Appendix F Proofs
F.1 Derivation of Target-Conditional Ordinary Least Squares
@@
  • The second term, $2\mathbb{E}\left[\sum_{m=1}^{M}Y^{\star}_{m}\theta^{\top}X^{\star}_{m}\Big|S^{\star}_{m}\right]$, can be rewritten using the linearity of expectation, the fact that $\theta$ is not random, and Assumption˜1:
@@
  • The third term, $\mathbb{E}\left[\sum_{m=1}^{M}(\theta^{\top}X^{\star}_{m})^{2}\Big|S^{\star}_{m}\right]$, is non-random by Assumption˜1 and can be rewritten as:
@@
as in Eq.˜4.
F.2 Proof of Theorem 7
In the main text, we stated our confidence interval for the 11NN choice of $\Psi$. We now state version for general non-negative matrices with columns summing to $1$.
@@ F.2 Proof of Theorem 7
Suppose $(S^{\star}_{m},X^{\star}_{m},Y^{\star}_{m})_{m=1}^{M}$ and $(S_{n},X_{n},Y_{n})_{n=1}^{N}$ satisfy Assumptions˜1, 4, 3 and 2. Define $w$ as in Algorithm˜1. For any $\Psi\in\mathbb{R}^{M\times N}$, a matrix with non-negative entries with columns summing to $1$, define $v^{\Psi}=w\Psi$. As in the main text, take $\hat{\theta}_{p}^{\Psi}=e_{p}^{\mathrm{T}}\left(X^{\star\mathrm{T}}X^{\star}\right)^{-1}X^{\star\mathrm{T}}\Psi Y$. Define $c=\sigma\|v^{\Psi}\|_{2}$, with $\sigma^{2}$ the variance of the additive noise from Assumption˜2. Define the (random) interval $I^{\Psi}$ as in Algorithm˜1 with known $\sigma^{2}$. Then with probability at least $1-\alpha$, $\theta^{\star}_{\text{OLS},p}\in I^{\Psi}$. That is, $I^{\Psi}$ has coverage (conditional on the test locations) at least $1-\alpha$.
@@
for $w:=e_{p}^{\mathrm{T}}\left(X^{\star\mathrm{T}}X^{\star}\right)^{-1}X^{\star\mathrm{T}}\in\mathbb{R}^{M}$ and $v^{\Psi}:=w\Psi\in\mathbb{R}^{N}$. By Assumption˜2, the righthand side of Eq.˜47 can be written
@@
We used the $\epsilon_{n}$ are independent and identically distributed with variance $\sigma^{2}$ (Assumption˜2).
@@
To bound the bias $b$, we use Assumption˜4 to write
@@
We can therefore apply Lemma˜6 to complete the proof. ∎
F.3 Computing an Upper Bound on the Bias of Our Estimand with Wasserstein-$1$ Distance
First of all observe that if $\sum_{m=1}^{M}w_{m}=1,\sum_{n=1}^{N}v_{n}^{\Psi}=1$, $v_{n}^{\Psi}\geq 0$ for $1\leq n\leq N$, and $w_{m}\geq 0$ for $1\leq m\leq M$ then the supremum in Eq.˜7 would be equal to a Wasserstein-$1$ distance by Kantorovich-Rubinstein duality [71, Theorem 5.10, Case 5.16].
Next, consider what happens if $\sum_{m=1}^{M}w_{m}-\sum_{n=1}^{N}v_{n}^{\Psi}\neq 0$. We show that the right-hand side of Eq.˜7 is infinite. Assume by contradiction that there exists a $C>0$ that upper bounds this supremum. Because $\sum_{m=1}^{M}w_{m}-\sum_{n=1}^{N}v_{n}^{\Psi}\neq 0$ and all constant functions are $L$-Lipschitz, for any $\gamma>0$, taking, for all $S$, $g(S)=G=\frac{C+\gamma}{\sum_{m=1}^{M}w_{m}-\sum_{n=1}^{N}v_{n}^{\Psi}}$